Baley:

A screw has a mechanical advantage of 12.5. How much force must you apply to the screw so that it has an output force of 1000 N

thesSMARTYpants:

First do cubs

thesSMARTYpants:

Circle the number Underline the keywords Bracket the question Eliminate the extra information Draw/Diagram

thesSMARTYpants:

Plz give me a rating if that helps

Vocaloid:

Moved to Physics

sillybilly123:

From Wiki [with emphasis added]: "Mechanical advantage is a measure of the force ***amplification*** achieved by using a tool, mechanical device or machine system." So, we can cast Mechanical advantage, which we can call \(\alpha\), as this: \(\alpha = \dfrac{F_{\text{out}}}{F_{\text{in}}}\) Then, with algebra, it becomes this: \(F_{\text{in}} = \dfrac{F_{\text{out}}}{\alpha} = \dfrac{1000}{12.5} \text{N}\) [This is a bad example, BTW, as screws are more about torque, but you should get the general idea behind leverage. ]